To do: Identify the oxidizing agent and reducing agent in the reactions.



1. 2Na2S2O3+I2→Na2S4O6+2NaI



2. 2FeCl3+H2S→2FeCl2+S+2HCl



3. 3Mg+N2→Mg3N24.AgCN+CN−→[Ag(CN)2]−

1
Expert's answer
2022-06-05T22:10:03-0400

1.

2Na2+S22+O32- +I20 = Na2+S42,5+O62- + 2Na+I-

4S2+ - 2e- => 4S2,5+ — reducing agent

I20 + 2e- => 2 I- — oxidizing agent

2.

2Fe3+Cl3- + H2+S2- = 2Fe2+Cl2- + S0 + 2H+Cl-

2Fe3+ +2e- => 2Fe2+ — oxidizing agent

2S- - 2e- => 2S0 — reducing agent

3.

3Mg0 + N20= Mg32+N23-

3Mg0 - 6e- => 3Mg2+ — reducing agent

N20 + 6e- => 2N3- — oxidizing agent

4.

Ag+(CN)- + (CN)- => [Ag+(CN)2-]-

Ag+ => Ag+

(CN)- => (CN)-

this is not a oxidation- reduction reaction